# Multi-step Inference over Unstructured Data

### v3

@@ Abstract
The advent of Large Language Models (LLMs) and Generative AI has revolutionized natural language applications across various domains. However, high-stakes decision-making tasks in fields such as medical, legal and finance require a level of precision, comprehensiveness, and logical consistency that pure LLM or Retrieval-Augmented-Generation (RAG) approaches often fail to deliver. At Elemental Cognition (EC), we have developed a neuro-symbolic AI platform to tackle these problems. The platform integrates fine-tuned LLMs for knowledge extraction and alignment with a robust symbolic reasoning engine for logical inference, planning and interactive constraint solving. We describe Cora, a Collaborative Research Assistant built on this platform, that is designed to perform complex research and discovery tasks in high-stakes domains. This paper discusses the multi-step inference challenges inherent in such domains, critiques the limitations of existing LLM-based methods, and demonstrates how Cora’s neuro-symbolic approach effectively addresses these issues. We provide an overview of the system architecture, key algorithms for knowledge extraction and formal reasoning, and present preliminary evaluation results that highlight Cora’s superior performance compared to well-known LLM and RAG baselines.
@@ 2.1 Life Science Research: Drug Discovery and Re-purposing
Figure 1 shows an example of using ChatGPT [^3] https://chat.openai.com/ (as of Apr 2024) for medical research. The example question is about exploring a potentially multi-hop relationship between Rheumatoid Arthritis (RA) and the inhibition of a particular kinase called IRAK4. As shown in the figure, there are four main classes of problems with this purely LLM-based approach: the inability to control the search or ranking process, since the LLM is a black-box that is not grounded to a specific corpus where one might apply filtering or ranking criteria; the inability to validate results without cross-checking the references (which defeats the purpose of an efficient research solution); hallucinated references which undermine credibility of the approach; and the ”needle-in-the-haystack” problem as valid answers that are not popular in the training data are unlikely to be surfaced by the LLM.
@@ 2.2 Macro-Economic Analysis: Multivariate Causal Inference
Consider the following macro-economic research problem: you are given a scenario that describes the state of a given economy, and the task is to understand the impact on a particular target concept. For example, If growth is falling in an Emerging Market country, and the country is facing high inflation, what is the likely impact on nominal bond yields?. As can be seen, the scenario contains multiple economic factors (falling economic growth, high inflation) that are present in a given context (Emerging Market country), and a good solution to this problem requires mapping out the relevant causal linkages between these factors and the target concept (nominal bond yields), and performing causal inference to net out the influences.
@@ 3.2 Knowledge Extraction using Statistical Models
For this problem, we have developed a transformer-based architecture called LUMEN that we use for concept typing, entity linking and relationship extraction, along with a fully automated domain-adaptation process that leverages our own fine-tuned LLMs to generate synthetic training data. The framework uses SLMs (Small Language Models such as Gemma-2B [13]) for increased throughput (sub-200ms latency per passage) without compromising on quality, by using high quality training data and novel contrastive loss functions. Details of this framework will be provided in a forthcoming technical publication.
@@ 3.3.2 Evidenced Graph Building and Symbolic Reasoning
As part of producing the map, the builder maps the specific relations (predicates) in the event structures stored in the Domain KB (described in Section 3.2) to the higher-order relationships such as ”influence” and ”trigger” in the Cogent Meta-model.
@@ 4 Preliminary Evaluation
We report results of a preliminary evaluation conducted in the medical domain. Further experiments are ongoing and we will share updated results and the dataset when this is completed.
@@ 4.1 Medical QA Eval
We collected 25 queries based on real questions from experts in the medical research domain. The queries were evaluated using four systems: GPT4-Turbo (a state-of-the-art LLM), Perplexity (RAG using web-search), Elicit (RAG using Semantic Scholar for doing scientific research) and Cora.
All four systems mentioned above were run on the questions and we asked each system to produce an answer with supporting/refuting evidence and cited sources. Since we do not have ground truth answers and explanations, we asked the domain experts to manually verify each of the systems’ results, and check that the generated answer/explanation justifies its claims, is both accurate and relevant, and that the sources it cites actually exist.
We have the following metrics:
@@
    Claim Density - average number of claims per answer - a measure of the quantity of information provided.
@@
    Citation Density - average number of real citations per claim - a measure of the amount of verification options.
@@
    Source Hallucination Rate - percentage of citations that are not valid (real) and scholarly sources - a measure of system hallucination.
@@
    Citation Rate - percentage of claims in the answer that are accompanied by real citations - a measure of verifiability of the answer.
@@
    Justification Rate - percentage of claims that are a correct paraphrase of a real citation - a measure of interpretation quality. Claims with non-existent sources are not justified as they are unverifiable. Since checking this requires manual effort, we imposed a max time-limit of 5 minutes on the domain expert to verify each claim.
@@
    Relevance Rate - percentage of claims that are justified and relevant to answering the question - a measure of relevance and quality of answer.
Results are shown in Table 1. Note that the metrics from 4-6 get progressively stricter, as a justified claim must also be cited, and a relevant claim must also be justified.
System Claim Density Citation Density Source Hallucination Rate Citation Rate Justi-fication Rate Relevance Rate
@@
Table 1: Complex QA Results with Evidence and cited Sources
@@ 5 Conclusions
We have described our approach and architecture in detail and presented experimental results validating its performance. Our results show that our solution delivers best in class performance for tackling multi-step causal inference problems on unstructured data when compared to pure LLM or state-of-the-art Retrieval Augmented Generation systems. We continue to evaluate our approach across a broader set of more complex problems and see promising results that we will share in the near future.

### v4

@@ Abstract
The advent of Large Language Models (LLMs) and Generative AI has revolutionized natural language applications across various domains. However, high-stakes decision-making tasks in fields such as medical, legal and finance require a level of precision, comprehensiveness, and logical consistency that pure LLM or Retrieval-Augmented-Generation (RAG) approaches often fail to deliver. At Elemental Cognition (EC), we have developed a neuro-symbolic AI platform to tackle these problems. The platform integrates fine-tuned LLMs for knowledge extraction and alignment with a robust symbolic reasoning engine for logical inference, planning and interactive constraint solving. We describe Cora, a Collaborative Research Assistant built on this platform, that is designed to perform complex research and discovery tasks in high-stakes domains. This paper discusses the multi-step inference challenges inherent in such domains, critiques the limitations of existing LLM-based methods, and demonstrates how Cora’s neuro-symbolic approach effectively addresses these issues. We provide an overview of the system architecture, key algorithms for knowledge extraction and formal reasoning, and present initial evaluation results that highlight Cora’s superior performance compared to well-known LLM and RAG baselines.
@@ 2.1 Life Science Research: Drug Discovery and Re-purposing
Figure 1 shows an example of using ChatGPT [^3] https://chat.openai.com/ (as of Apr 2024) for medical research. The example question is about exploring a potentially multi-hop relationship between Rheumatoid Arthritis (RA) and the inhibition of a particular kinase called IRAK4. As shown in the figure, there are four main classes of problems with this purely LLM-based approach: the inability to control the search or ranking process, since the LLM is a black-box that is not grounded to a specific corpus where one might apply filtering or ranking criteria; the inability to validate results without cross-checking the references (which defeats the purpose of an efficient research solution); hallucinated references which undermine credibility of the approach; and the “needle-in-the-haystack” problem as valid answers that are not popular in the training data are unlikely to be surfaced by the LLM.
@@ 2.2 Macro-Economic Analysis: Multivariate Causal Inference
Consider the following macro-economic research problem: you are given a scenario that describes the state of a given economy, and the task is to understand the impact on a particular target concept. For example, “If growth is falling in an Emerging Market country, and the country is facing high inflation, what is the likely impact on nominal bond yields?”. As can be seen, the scenario contains multiple economic factors (“falling economic growth”, “high inflation”) that are present in a given context (“Emerging Market country”), and a good solution to this problem requires mapping out the relevant causal linkages between these factors and the target concept (“nominal bond yields”), and performing causal inference to net out the influences.
@@ 3.2 Knowledge Extraction using Statistical Models
For this problem, we have developed a transformer-based architecture for Linking and Understanding Mentions using Encoder Networks, LUMEN, that we use for concept typing, entity linking and relationship extraction, along with a fully automated domain-adaptation process that leverages our own fine-tuned LLMs to generate synthetic training data. The framework uses SLMs (Small Language Models such as Gemma-2B [13]) for increased throughput (sub-200ms latency per passage) without compromising on quality, by using high quality training data and novel contrastive loss functions. Details of this framework will be provided in a forthcoming technical publication.
@@ 3.3.2 Evidenced Graph Building and Symbolic Reasoning
As part of producing the map, the builder maps the specific relations (predicates) in the event structures stored in the Domain KB (described in Section 3.2) to the higher-order relationships such as “influence” and “trigger” in the Cogent Meta-model.
@@ 4 Initial Evaluation
We report results of an initial evaluation conducted in the medical domain.
@@ 4.1 Medical QA Eval
We collected two sets of 25 queries each based on real questions from experts in the medical research domain.
  • Representative Queries: The first set is composed of queries that are representative of the types of queries experts could ask of a medical QA system. This set kept the queries as they were asked by the experts. (see Appendix 7.1 for the list of the queries.)
  • Multi-hop Queries: The second set was specifically selected to evaluate “multi-hop” question answering. A satisfactory answer to such queries would need to outline complex causal relations involving multiple intermediary causal factors. In order to avoid confounds due to surface form variations (which is already addressed by the first data set) and to facilitate the evaluation of answer complexity, the queries were reformatted to follow the format “How does X impact Y?”. (see Appendix 7.2 for the list of the queries.)
The queries were evaluated using four systems:
  • GPT4-Turbo [^6] https://platform.openai.com/playground/chat?models=gpt-4-turbo: a state-of-the-art LLM
  • Perplexity [^7] https://www.perplexity.ai/: RAG using web-search
  • Elicit [^8] https://elicit.com/: RAG using Semantic Scholar for doing scientific research
  • Cora: Our solution. [^9] Although our systems allows a user to further refine an answer, we here evaluate the direct answer from Cora to a user’s question without any further user interactions.
All four systems mentioned above were run on the questions and we asked each system to produce an answer with supporting/refuting evidence and cited sources.
We focus on evaluation on intrinsic task of verification and relevance of answers as opposed to direct comparison with what would be a necessarily incomplete gold standard answer as each question could have multiple equally correct answers. Domain experts manually reviewed each of the systems’ results. They were asked to check that the generated answer/explanation justifies its claims, is both accurate and relevant, and that the sources it cites actually exist.
We designed the following metrics to assess whether or not a user can verify the claims presented and whether or not those are relevant:
@@
    Claim Density: Average number of claims per answer. A measure of the quantity of information provided.
@@
    Citation Density: Average number of real citations per claim. A measure of the amount of verification options.
@@
    Source Hallucination Rate: Percentage of citations that are not valid (real) and scholarly sources. A measure of system hallucination.
@@
    Citation Rate: Percentage of claims in the answer that are accompanied by real citations. A measure of verifiability of the overall answer.
@@
    Justification Rate: Percentage of claims that are a correct paraphrase of a real citation. A measure of interpretation quality. Claims with non-existent sources are not justified as they are unverifiable. Since checking this requires manual effort, we imposed a max time-limit of 5 minutes on the domain expert to verify each claim.
@@
    Relevance Rate. Percentage of claims that are justified and relevant to answering the question. A measure of relevance and quality of answer.
Results are shown in Table 1 and 2. Note that the metrics from 4-6 get progressively stricter, as a justified claim must also be cited, and a relevant claim must also be justified.
System Claim Density Citation Density Source Hallucination Rate Citation Rate Justification Rate Relevance Rate
@@
Table 1: Representative Queries. Results with Evidence and cited Sources.
System Claim Density Citation Density Source Hallucination Rate Citation Rate Justification Rate Relevance Rate
GPT4-Turbo 4.16 (+0.64) 1.01 (+0.38) 31.4% (-11.46) 64.42% (+16.69) 27.88% (+2.88) 22.12% (-1.74)
Perplexity 4.76 (+1.16) 0.59 (+0.21) 0.01% (+0.01) 32.77% (+10.55) 17.65% (-3.46) 11.76% (+1.76)
Elicit 5.00 (+1.04) 1.36 (-0.17) 0.01% (-3.81) 98.40% (+2.44) 86.40% (+3.57) 60.80% (-8.90)
Cora 5.36 (+0.32) 1.14 (-0.12) 0.00% (+0.00) 98.51% (-1.49) 90.30% (-3.35) 86.57% (+1.65)
Table 2: Multi-hop Queries. Results with Evidence and cited Sources. Values in parentheses indicate the difference with Representative Queries results of Table 1
In addition, for the answers to the Multi-hop Queries dataset, we were interested in quantifying the complexity of the answers generated. Given that the queries are of the type “What is the impact of X on Y?”, a domain expert recorded for each answer:
  1. 1.
    Maximum Number of Hops: Maximum number of hops (relations) tying the source (X) to the target (Y) in a reasoning chain. (e.g. ’X causes A1 causes A2 causes Y’ would be a 3-hop reasoning chain since it is composed of 3 causal relations). This an indication of the depth of the answer.
  2. 2.
    Number of Concepts: Number of concepts presented in the answer that are directly relevant to the explaining the mechanism. This is an indication of coverage of the answer.
Results are shown in Table 3.
System Maximum Number of Hops Number of Concepts
[…]
Table 3: Multi-hop Queries. Answer Complexity Analysis.
@@ 4.2.1 Results on Representative Queries data
4.2.2 Results on Multi-hop Queries data
Results for the Multi-hop Queries are shown in Table 2. All the points discussed above still hold in the case of multi-hop queries. GPT-4 Turbo and Perplexity, systems that had the lowest performances in the Representative Queries, display similar poor results on multi-hop queries. Crucially, looking at the final relevance rate, there is no drop in quality of answer in Cora while Elicit’s final relevance score on the other hands drops by almost nine points.
The answer complexity analysis shown in Table 3 add another dimension to the results. A pure LLM solution such as GPT-4 Turbo generates answers with a relatively high number of concepts (good coverage) and with the longest reasoning chains of more than 2 hops on average (good depth). However, as shown in Table 2, most of its claims are not justified and irrelevant. Conversely, Elicit, a system tailored for research, has a higher rate of justification and relevance but its answers contain fewer concepts (low coverage) and with fewer hops (lack of depth). Cora’s answers combine both a high coverage and depth with justified and relevant claims.
Cora distinguishes itself by the stability of its performances as well as the depth and coverage of its answers. This is due to EC’s implementation that integrates LLMs with symbolic causal models into the answer generation process. This allows for the iterative exploration, capture, verification and refinement of relevant knowledge extracted from unstructured domain data into an integrated causal graph. This graph in turn provides a solid semantic scaffold to guide the text generation process.
@@ 5 Conclusions
We have described our approach and architecture in detail and presented experimental results validating its performance. Our results show that our solution delivers best in class performance at answering complex research questions requiring multi-step exploration, reasoning and justification over a knowledge-intensive domain.